OpenStudy (anonymous):

Simplify. -4.5[2x - (0.5 -1.4)] A. -9x + 0.9 B. -9x - 0.9 C. -9x - 4.05 D. 9x - 4.05

OpenStudy (anonymous):

@Daniellelovee @dtan5457

OpenStudy (daniellelovee):

distributive property

OpenStudy (ajspeller):

Begin in the innermost set of parenthesis.

OpenStudy (ajspeller):

Order of operations: PEMDAS

OpenStudy (anonymous):

0.5 -1.4 = 0.9 - 2x = -1.1 * -4.5 = 4.95

OpenStudy (daniellelovee):

not quite...

OpenStudy (anonymous):

Help a little more plz

OpenStudy (daniellelovee):

i will hold on I'll show you how I would solve it

OpenStudy (anonymous):

okay

OpenStudy (ajspeller):

 I will leave the rest to you.